A 5.82 g silver coin is dissolved in nitric acid. When sodium chloride is added to the solution, all the silver is precipitated as AgCl. The AgCl precipitate weights 7.20 g. What is the percentage of solver in coin?


(A)        98%


(B)        93.1%


(C)        86%


 


(D)        82%

Grade:upto college level

1 Answers

Shane Macguire
10 years ago

Fraction of Ag in AgCl = 107.9/143.3= 0.753  weight of Ag in 7.20 g AgCl

 

                = 0.753 × 7.20 g = 5.42 g Ag

 

        Hence the 5.82 g coin contains 5.42 g Ag = 0

 

        Fraction of Ag in coin = 5.42 g/5.82 g

 

                = 0.931 = 93.1 % Ag